OpenStudy (anonymous):

Will give medal. c: Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 Equation 2 |5x + 6| = 41 |2x + 13| = 28

OpenStudy (aum):

Go ahead and solve each equation separately. When using absolute bars, you can get rid of it as follows: If |A| = B it means either A = B or A = -B |5x + 6| = 41 means: 5x + 6 = 41 or 5x + 6 = -41 solve each separately.

OpenStudy (anonymous):

5x + 6 = 41 - 6 - 6 5x= 35 5x/5 = 35/5 x=7 5x + 6 = -41 - 6 - 6 5x = -35 5x/5= -35/5 x= -7 is this right?

OpenStudy (texaschic101):

The first one is correct... The 2nd one is set up like this... -(5x + 6) = 41 -5x - 6 = 41 -5x = 41 + 6 -5x = 47 x = -47/5 ============= | 2x + 13 | = 28 2x + 13 = 28 2x = 28 - 13 2x = 15 x = 7.5 -(2x + 13) = 28 -2x - 13 = 28 -2x = 28 + 13 -2x = 41 x = - 20.5 ================= read this... http://www.purplemath.com/modules/solveabs.htm

OpenStudy (anonymous):

ohh, my options are worded though

OpenStudy (anonymous):

i'll post them now

OpenStudy (anonymous):

Eq. 1 has more solutions than Eq. 2 Eq1 and Eq2 have the same number of solutions Eq2 has more solutions than Eq1 The number of solutions cannot be determined

OpenStudy (anonymous):

@texaschic101 which would it be? the options is the part that confused me

OpenStudy (radar):

Look at what you got and what texaschic101 got. Note that equation had two solutions and only two. Equation also had 2 and only 2 solutions. which option fits the situation?

OpenStudy (radar):

Each equation had 2 and only 2 solutions.